OpenStudy (hatahateo):

Help, I'll fan and medal Write a linear factorization of the function. f(x) = x4 + 36x2

OpenStudy (welshfella):

first ind the GCF of the 3 terms (greatest common factor). That is the greatest term which divides exactly into both terms.

OpenStudy (welshfella):

* GCF of the 2 terms

OpenStudy (hatahateo):

I think it is f(x) = x2(x + 6i)(x - 6i)

OpenStudy (welshfella):

yes thats correct

OpenStudy (welshfella):

so you have studied complex numbers ?

OpenStudy (hatahateo):

Yes, a little bit. Can you help with another?

OpenStudy (welshfella):

ok

OpenStudy (hatahateo):

Do you want me to open a new question?

OpenStudy (hatahateo):

State how many imaginary and real zeros the function has. f(x) = x3 + 5x2 - 28x - 32

OpenStudy (hatahateo):

I said it has 0 imaginary and 3 real numbers

OpenStudy (welshfella):

oh - you have used Descarte's Rule of Signs right?

OpenStudy (hatahateo):

Yes

OpenStudy (welshfella):

I cant remember the details of that. it will be either 1 real and 2 imaginary or 3 real as you said.

OpenStudy (hatahateo):

So, I am wrong with the 0 imaginary and 3 real?

OpenStudy (welshfella):

i see that f(-1) = 0 so x = -1 is one root

OpenStudy (welshfella):

No - you might be right

OpenStudy (hatahateo):

Alright

OpenStudy (welshfella):

i see that the last number is -32 so see if +/-8 or +/- is a root. ( applying the rational root theorem)

OpenStudy (welshfella):

f(4) = 4^3 + 5(4)^2 - 28(4) - 32

OpenStudy (hatahateo):

Okay

OpenStudy (welshfella):

= 64 + 80 - 112 - 32 = ?

OpenStudy (hatahateo):

Hang on

OpenStudy (hatahateo):

0

OpenStudy (robtobey2):

\[x^3+5x^2-28x-32=(x-4) (x+1) (x+8) \]

OpenStudy (welshfella):

looks like I got lucky... yes 4 is a root so ur answer is correct

OpenStudy (hatahateo):

Thanks! I have another one too

OpenStudy (welshfella):

pleass repost any more

OpenStudy (hatahateo):

???

OpenStudy (welshfella):

open a new question

OpenStudy (mww):

yes rational root theorem is nice for some of those polys